# Radiative-Structured Neural Operator for Continuous and Extrapolative Spectral Super-Resolution

### 07 Mar 2026 v2

@@ Abstract
Spectral super-resolution (SSR) aims to reconstruct hyperspectral images (HSIs) from multispectral observations, with broad applications in computer vision and remote sensing. Deep learning-based methods have been widely used, but they often treat spectra as discrete vectors learned from data, rather than continuous curves constrained by physics principles, leading to unrealistic predictions and limited applicability. To address this challenge, we propose the Radiative-Structured Neural Operator (RSNO), which learns a continuous mapping for spectral super-resolution while enforcing physical consistency under the radiative prior. The proposed RSNO consists of three stages: upsampling, reconstruction, and refinement. In the upsampling stage, we leverage prior information to expand the input multispectral image, producing a physically plausible hyperspectral estimate. Subsequently, we adopt a neural operator backbone in the reconstruction stage to learn a continuous mapping across the spectral domain. Finally, the refinement stage imposes a hard constraint on the output HSI to eliminate color distortion. The upsampling and refinement stages are implemented via the proposed angular-consistent projection (ACP), which is derived from a non-convex optimization problem. Moreover, we theoretically demonstrated the optimality of ACP by null-space decomposition. Various experiments validate the effectiveness of the proposed approach across conventional spectral super-resolution, continuous spectral reconstruction, and infrared extrapolation.
@@ I Introduction
Traditional SSR approaches mainly include early model-based methods [29], regression-based methods [33, 44], and sparse representation methods [1, 49]. These methods primarily rely on prior knowledge and physical constraints, which contribute to robustness and interpretability. However, they often fail to fully leverage large-scale data, which may lead to underfitting.
Advances in deep learning have led to a range of approaches based on deep neural networks. Numerous methods employing convolutional neural networks (CNNs) [38, 30] and attention mechanisms [3, 45, 20] lead to notable improvements and outperform traditional approaches. Deep learning-based methods can also be categorized according to their problem settings. For example, fusion-based methods [34, 10] use paired low-resolution HSIs and high-resolution MSIs to extract and fuse spectral-spatial information, producing high-resolution hyperspectral images. However, the practical application of fusion-based methods is hindered by the difficulty of obtaining registered HSI-MSI pairs. Thus, single-image spectral super-resolution methods [41, 26, 20, 48, 51] have gained attention, as they require only the input MSI or RGB image at the inference stage. For example, Zhou et al. [51] proposed a CNN-Transformer hybrid network for directly reconstructing hyperspectral images from RGB inputs, thereby enhancing spectral super-resolution. Unsupervised spectral super-resolution has also been extensively studied to address the challenge of acquiring paired MSI-HSI data [35, 8].
@@
To address this challenge, we propose the Radiative-Structured Neural Operator (RSNO), which models spectra in infinite-dimensional space and incorporates physically grounded spectral priors derived from atmospheric radiative transfer models. Inspired by the two-stage paradigm [38, 30, 21], the proposed framework comprises two standard stages, upsampling and reconstruction, and an additional refinement stage, as illustrated in Figure 1. In the upsampling stage, the spectral response function (SRF) and radiative transfer priors are leveraged to generate an image with higher spectral resolution. With the upsampled HSI, we employ a data-driven neural operator backbone to reconstruct a finer HSI estimate. Since the neural operator is resolution-invariant and coordinate-based [22, 18], we can compute the estimated HSI at arbitrary spectral scales. Finally, we refine the HSI by projecting it onto the affine space defined by the SRF, thereby ensuring consistency with the original MSI after degradation. The upsampling and refinement stages employ our proposed angular-consistent projection (ACP), a novel methodology derived from a non-convex affine space projection problem based on cosine similarity. It not only leverages physical priors to upsample the MSI but also enforces hard constraints to guarantee reconstruction consistency. Additionally, the neural operator backbone adopts a U-shaped architecture equipped with spectral-aware convolution (SAC) layers to capture multi-scale spatial–spectral features. Experiments on real-world datasets show that RSNO not only performs competitively in conventional finite-dimensional SSR, but also enables continuous spectral reconstruction and infrared extrapolation, highlighting its effectiveness and flexibility in practical applications.
@@
  • We formulate spectral super-resolution as a continuous regression problem in infinite-dimensional space, supporting both continuous interpolation and infrared extrapolation;
@@ II-A Continuous Spectral Super-Resolution
Continuous spectral super-resolution aims at recovering spectral curves at arbitrary resolutions. Xu et al. [45] first introduced a framework that reconstructs spectral images with an arbitrary number of bands and demonstrated superior performance over discrete methods in fixed-band reconstruction tasks. Beyond this, they further upgraded the neural attention mapping module in their subsequent work [46], thereby enabling continuous reconstruction in both the spatial and spectral domains with state-of-the-art performance among discrete baselines. For HSI reconstruction of the Coded Aperture Snapshot Spectral Imaging (CASSI) system, Chen et al. [5] proposed a spectral-wise attention mechanism with structural designs to achieve continuous HSI recovery. Similarly, Li et al. [19] proposed a lightweight architecture that supports continuous reconstruction across both spatial and spectral dimensions within the CASSI framework.
These approaches are all built upon implicit neural representations (INRs) [6]. Originating from the neural radiance field (NeRF) [31], INR implicitly represents natural and complex images continuously and has inspired a wide range of arbitrary-scale super-resolution methods in computer vision [32].
@@ II-B Physics-Informed Spectral Super-Resolution
While data-driven methods have achieved strong performance in spectral super-resolution, their lack of explicit physical constraints has motivated the development of physics-informed approaches [23, 13, 25, 16, 43, 7]. These methods seek to integrate physical models or domain-specific priors into learning-based frameworks to enhance reconstruction fidelity and interpretability.
Existing physics-informed SSR methods can be broadly categorized into regularization-based and model-embedded approaches, depending on how physical priors are incorporated. Regularization-based methods typically incorporate physical knowledge at the objective level in the form of auxiliary losses or regularization terms. For example, Wu et al. [43] embedded comprehensive semantic priors to regularize and optimize the solution space of SSR. Model-embedded methods integrate physical knowledge directly into the model design, where physical structures are reflected in the network architecture, parameterization, or computational modules. For example, Liu et al. [25] embedded a spectral unmixing model into the reconstruction framework and explicitly accounted for solar radiation and atmospheric absorption effects. In addition, several works [23, 16, 8] approached SSR from the perspective of a linear inverse problem and derived null-space priors to constrain the solution space, realizing high-fidelity reconstruction.
@@ II-C Operator Learning
Neural operators have a wide range of mathematical and scientific applications. Beyond their utility in scientific computing, neural operators have also been applied to super-resolution tasks in computer vision and remote sensing [42, 50, 14]. Their inherent resolution-invariance property enables continuous super-resolution, making them particularly well-suited for these applications.
@@ III Methodology
Figure 1: Overall framework of the proposed method. The input MSI undergoes three main stages to output the final HSI $\hat{X}$: (1) Upsampling, where we expand the input MSI to the upsampled HSI, (2) Reconstruction, where we employ the neural operator to transform the upsampled HSI to a finer HSI estimate, and (3) Refinement, where we impose a hard constraint on the final result to enforce zero reconstruction error and eliminate color distortion. The upsampling and refinement stages both employ the ACP method, but with different objectives. In the upsampling stage, ACP incorporates radiative spectral prior to produce a radiative-structured and physically plausible HSI estimate, whereas in the refinement stage, it projects the HSI onto the affine space defined by (4).
@@ IV Experiments
Figure 3: Visual comparison of pseudo-RGB reconstruction results across different sites. Each row presents the pseudo-RGB images of one spectral reconstruction method applied to one scene from six different sites, while each column compares the pseudo-RGB images of various methods on the same scene within a single site. The MRAE reconstruction error is also calculated and presented on top of the image.
In this section, we present numerical experiment results to validate our approach under various settings. We begin by introducing the datasets in Section IV-A and outlining the basic experimental setups in Section IV-B. We first evaluate our model under the conventional discrete SSR setting, with results reported in Section IV-C. We then assess its performance in continuous spectral super-resolution and infrared spectral extrapolation scenarios in Section IV-D. Finally, an ablation study is presented in Section IV-E to analyze how and why the proposed design is effective.
Figure 4: Visual comparison of MSE error maps across different sites. Each row presents the error maps of a specific spectral reconstruction method on a scene from six sites, while each column shows the error maps of different methods applied to the same scene of a given site.
(a) Blodgett Forest
(b) Cuprite
(c) Everglades
(d) Los Angeles
(e) Madison
(f) Yellowstone
@@ IV-C Discrete Spectral Super-Resolution
RSNO (Ours) 4.8 0.160 45.400 0.137 0.973
We conduct comparative experiments in conventional discrete SSR setting to evaluate the performance of the proposed RSNO against several representative methods, including UnGUN [35], INR [48], NeSR [45], and NeSSR [46]. All models are evaluated under the same training-testing protocol.
Table II reports the overall quantitative results on the test datasets. Among all baselines, our RSNO consistently achieves the best performance, with the lowest MRAE and SAM, the highest PSNR and SSIM, and the fewest parameters among high-performing methods. These results highlight the efficiency and effectiveness of our design. Furthermore, we also conduct per-site evaluations on six representative scenes, and the results are summarized in Table III. RSNO consistently delivers top or near-top performance in most scenarios, particularly excelling in complex scenes such as rural areas. This demonstrates that the model is highly generalizable and adaptable to heterogeneous landforms and spectral distributions.
We also conduct a visual comparison among all models, as shown in Figure 3, Figure 4, and Figure 5. Figure 3 demonstrates that our method is capable of faithfully restoring RGB images, whereas other methods exhibit visual artifacts or unnatural color distortions. This advantage mainly stems from the refinement step, which explicitly enforces zero reconstruction loss with ACP, ensuring accurate color consistency. The error maps in Figure 4 further highlight the superiority of our method. Additionally, we plot the predicted spectrum on six sites, which are shown in Figure 5. Our method yields more natural and physically plausible spectra that better align with the ground-truth, especially in wavelength regions affected by atmospheric absorption, whereas other approaches tend to either overfit noisy variations or underfit the true reflectance shape.
In general, both quantitative and qualitative comparisons indicate that RSNO achieves state-of-the-art performance and more stable results in discrete setting across varied geographical contexts with fewer parameters.
@@
RSNO (Ours) 43.296 0.112 53.423 0.049 42.188 0.183 41.884 0.148 41.601 0.125 45.575 0.221
TABLE IV: Quantitative Evaluation of Continuous and Extrapolative Spectral Reconstruction
Method Continuous Interpolation Infrared Extrapolation
@@
NeSR [45] 0.565 34.116 0.374 0.681 0.967 30.709 0.860 0.532
NeSSR [46] 0.563 34.309 0.528 0.685 1.054 30.049 0.664 0.584
RSNO (Ours) 0.220 43.144 0.194 0.944 0.707 34.457 0.641 0.765
IV-D Continuous and Extrapolative Spectral Super-Resolution
To validate the effectiveness of our continuous modeling and evaluate the generalization and flexibility of our method, we conduct experiments on continuous spectral reconstruction and infrared spectral extrapolation.
In the continuous reconstruction setting, we train the model using only a small number of uniformly downsampled spectral bands and test it on the full spectral resolution. This simulates an interpolation task in the spectral domain, where the model needs to infer missing spectral information between known bands. In the infrared extrapolation setting, we train the model using only the visible spectrum range from 400 nm to 1000 nm, but test the model on the complete spectral range from 400 nm to 2500nm, which evaluates the ability to reconstruct the entire hyperspectral signal, including the never-seen infrared region. This corresponds to a spectral extrapolation task, which is highly challenging due to the lack of direct observation in the infrared spectral range.
In our experiments, we apply $2\times$ downsampling for continuous spectral reconstruction and restrict the spectrum to wavelengths below 1000 nm for extrapolative reconstruction. We compare our method against two other continuous methods: NeSR and NeSSR. Table IV summarizes the quantitative results. Our method achieves best performance in both settings, indicating strong generalization to unseen or sparsely-sampled spectral distributions.
The good performance of RSNO stems from two factors: the neural operator backbone ensuring resolution-invariant spectral interpolation, and the ACP method guided by radiative spectral prior ensuring radiative-structured extrapolation from visible to infrared bands. These results highlight RSNO’s robustness and generalization ability, showing that it can both interpolate missing spectra and extrapolate beyond observed bands, making it suitable for practical hyperspectral applications.
@@ IV-E Ablation Study
where $\alpha$ is the trade-off parameter, and we set $\alpha=0.5$ in the ablation experiment. The results are listed in Table V.
The results indicate that both the radiative prior and the final refinement play critical roles in the overall performance. Removing the radiative prior in the upsampling stage reduces PSNR and increases SAM, showing that physically informed guidance is crucial for producing an initial estimate close to the target. However, eliminating the refinement stage and replacing it with soft regularization degrades both PSNR and SAM more significantly, suggesting that hard constraints in the output space are more effective than radiative prior in preserving spectral fidelity. Overall, these findings highlight that the ACP method plays a crucial role in ensuring accurate and physically consistent spectral reconstruction.
@@ V Conclusion
In this article, we propose a novel physics-informed approach for continuous and extrapolative spectral reconstruction. By continuously reformulating the problem and embedding physical priors within a data-driven framework, our method yields more realistic reconstructions and effectively balances interpretability and adaptability. Experiments across diverse real-world scenes and problem settings indicate better reconstruction accuracy and robustness, especially in the continuous setting. This work highlights the potential of continuous modeling of spectra and integrating domain knowledge with deep learning to advance spectral super-resolution in practical applications.

### 12 Mar 2026 v3

# Radiative-Structured Neural Operator for Continuous Spectral Super-Resolution

@@ Abstract
Spectral super-resolution (SSR) aims to reconstruct hyperspectral images (HSIs) from multispectral observations, with broad applications in computer vision and remote sensing. Deep learning-based methods have been widely used, but they often treat spectra as discrete vectors learned from data, rather than continuous curves constrained by physics principles, leading to unrealistic predictions and limited applicability. To address this challenge, we propose the Radiative-Structured Neural Operator (RSNO), which learns a continuous mapping for spectral super-resolution while enforcing physical consistency under the radiative prior. The proposed RSNO consists of three stages: upsampling, reconstruction, and refinement. In the upsampling stage, we leverage prior information to expand the input multispectral image, producing a physically plausible hyperspectral estimate. Subsequently, we adopt a neural operator backbone in the reconstruction stage to learn a continuous mapping across the spectral domain. Finally, the refinement stage imposes a hard constraint on the output HSI to eliminate color distortion. The upsampling and refinement stages are implemented via the proposed angular-consistent projection (ACP), which is derived from a non-convex optimization problem. Moreover, we theoretically demonstrated the optimality of ACP by null-space decomposition. Various experiments validate the effectiveness of the proposed approach in both discrete and continuous spectral super-resolution.
@@ I Introduction
Traditional SSR approaches mainly include model-based methods [29], regression-based methods [32, 43], and sparse representation methods [1, 48]. These methods primarily rely on prior knowledge and physical constraints, which contribute to robustness and interpretability. However, they often fail to fully leverage large-scale data, which may lead to underfitting.
Advances in deep learning have led to a range of approaches based on deep neural networks. Numerous methods employing convolutional neural networks (CNNs) [37, 30] and attention mechanisms [3, 44, 20] lead to notable improvements and outperform traditional approaches. Deep learning-based methods can also be categorized according to their problem settings. For example, fusion-based methods [33, 10, 53] use paired low-resolution HSIs and high-resolution MSIs to extract and fuse spectral-spatial information, producing high-resolution hyperspectral images. However, the practical application of fusion-based methods is hindered by the difficulty of obtaining registered HSI-MSI pairs. Thus, single-image spectral super-resolution methods [40, 26, 20, 47, 51] have gained attention, as they require only the input MSI or RGB image at the inference stage. Unsupervised spectral super-resolution has also been extensively studied to address the challenge of acquiring paired MSI-HSI data [34, 8].
@@
To address this challenge, we propose the Radiative-Structured Neural Operator (RSNO), which models spectra in infinite-dimensional space and incorporates physically grounded spectral priors derived from atmospheric radiative transfer models. Inspired by the two-stage paradigm [37, 30, 21], the proposed framework comprises two standard stages, upsampling and reconstruction, and an additional refinement stage, as illustrated in Figure 1. In the upsampling stage, the spectral response function (SRF) and radiative transfer priors are leveraged to generate an image with higher spectral resolution. With the upsampled HSI, we employ a data-driven neural operator backbone to reconstruct a finer HSI estimate. Since the neural operator is resolution-invariant and coordinate-based [22, 18], we can compute the estimated HSI at arbitrary spectral scales. Finally, we refine the HSI by projecting it onto the affine space defined by the SRF, thereby ensuring consistency with the original MSI after degradation. The upsampling and refinement stages employ our proposed angular-consistent projection (ACP), a novel methodology derived from a non-convex affine space projection problem based on cosine similarity. It not only leverages physical priors to upsample the MSI but also enforces hard constraints to guarantee reconstruction consistency. Additionally, the neural operator backbone adopts a U-shaped architecture equipped with spectral-aware convolution (SAC) layers to capture multi-scale spatial–spectral features. Experiments on real-world datasets show that RSNO not only performs competitively in conventional finite-dimensional SSR, but also enables continuous spectral reconstruction, highlighting its effectiveness and flexibility in practical applications.
@@
  • We formulate spectral super-resolution as a continuous regression problem in infinite-dimensional space, supporting arbitrary-scale spectral super-resolution;
@@ II-A Continuous Spectral Super-Resolution
Continuous spectral super-resolution aims at recovering spectral curves at arbitrary resolutions. Xu et al. [44] first introduced a framework that reconstructs spectral images with an arbitrary number of bands and demonstrated superior performance over discrete methods in fixed-band reconstruction tasks. Beyond this, they further upgraded the neural attention mapping module in their subsequent work [45], thereby enabling continuous reconstruction in both the spatial and spectral domains with state-of-the-art performance among discrete baselines. Besides, Chen et al. [5] and Li et al. [19] studied continuous HSI reconstruction within the Coded Aperture Snapshot Spectral Imaging (CASSI) framework. Additionally, Zhou et al. [52] proposed an efficient method for arbitrary scale spatial-spectral super-resolution based on functional tensor decomposition.
Many continuous approaches are built upon implicit neural representations (INRs) [6], which implicitly represent natural and complex images continuously and have inspired a wide range of arbitrary-scale super-resolution methods in computer vision [50, 31].
@@ II-B Physics-Informed Spectral Super-Resolution
While data-driven methods have achieved strong performance in spectral super-resolution, their lack of explicit physical constraints has motivated the development of physics-informed approaches [23, 13, 25, 16, 42, 7]. These methods seek to integrate physical models or domain-specific priors into learning-based frameworks to enhance reconstruction fidelity and interpretability. For example, Wu et al. [42] incorporated comprehensive semantic priors to regularize and optimize the solution space of SSR; Liu et al. [25] embedded a spectral unmixing model into the reconstruction framework and explicitly accounted for solar radiation and atmospheric absorption effects. In addition, several works [23, 16, 8] approached SSR from the perspective of a linear inverse problem and derived null-space priors to constrain the solution space, realizing high-fidelity reconstruction.
@@ II-C Operator Learning
Neural operators have a wide range of mathematical and scientific applications. Besides, neural operators have also been applied to super-resolution tasks in computer vision and remote sensing [41, 49, 14, 53]. Their inherent resolution-invariance property enables continuous super-resolution, making them particularly well-suited for these applications.
@@ III Methodology
Figure 1: Overall framework of the proposed method. The input MSI undergoes three main stages to output the final HSI $\hat{X}$: (1) Upsampling, where we expand the input MSI to the upsampled HSI, (2) Reconstruction, where we employ the neural operator to transform the upsampled HSI to a finer HSI estimate, and (3) Refinement, where we impose a hard constraint on the final result to enforce zero reconstruction error and eliminate color distortion.
@@ IV Experiments
Figure 3: Visual comparison of reconstruction results across different sites. Each row corresponds to a spectral reconstruction method, while each column corresponds to a scene from a specific site. In (b), the MRAE reconstruction error is shown on top of each image.
In this section, we present numerical experiment results to validate our approach under various settings. We begin by introducing the datasets in Section IV-A and outlining the basic experimental setups in Section IV-B. We first evaluate our model under the conventional discrete SSR setting, with results reported in Section IV-C. We then assess its performance in continuous spectral super-resolution in Section IV-D. Finally, an ablation study is presented in Section IV-E to analyze how and why the proposed design is effective.
@@ IV-C Discrete Spectral Super-Resolution
(a) Blodgett Forest
(b) Cuprite
(c) Everglades
(d) Los Angeles
(e) Madison
(f) Yellowstone
We conduct comparative experiments in the conventional discrete SSR setting to evaluate the performance of the proposed RSNO against several representative methods, including UnGUN [34], INR [47], NeSR [44], and NeSSR [45]. All models are evaluated under the same training-testing protocol.
Table II reports the overall quantitative results on the test datasets. Among all baselines, our RSNO consistently achieves the best performance, with the lowest MRAE and SAM, the highest PSNR and SSIM, and the fewest parameters among high-performing methods. These results highlight the efficiency and effectiveness of our design. Besides, we conduct per-site evaluations to further compare the performance of the four methods. The results are summarized in Table III. RSNO consistently delivers top or near-top performance in most scenarios, particularly excelling in complex scenes such as desert and wetland. However, results in the urban area and mountainous area show that RSNO achieves slightly inferior performance compared with the competing methods.
We also conduct a visual comparison among all models, as shown in Figure 3. From error maps we can observe that RSNO achieves lower MSE in most scenarios, except for Los Angeles, which is the most complex scene with dense man-made structures. Pseudo-RGB reconstruction results demonstrate that our method is capable of faithfully restoring RGB images, whereas other methods exhibit visual artifacts or unnatural color distortions. This advantage mainly stems from the refinement step, which explicitly enforces zero reconstruction loss with ACP, ensuring accurate color consistency. Additionally, we plot the predicted spectrum on six sites, which are shown in Figure 4. Our method yields more natural and physically plausible spectra that better align with the ground truth, especially in wavelength regions affected by atmospheric absorption. In contrast, other approaches tend to either overfit noisy variations or underfit the true reflectance shape.
In general, both quantitative and qualitative comparisons indicate that RSNO achieves state-of-the-art performance and more stable results in the discrete setting across varied geographical contexts with fewer parameters. However, certain limitations are also observed. In particular, RSNO may suffer from less accurate predictions in complex scenarios such as the urban area, where highly intricate man-made structures pose significant challenges. Addressing this issue remains an important direction for future work.
@@
[…]
@@
[…]
TABLE IV: Quantitative Evaluation of Continuous Spectral Reconstruction
Method $2\times$ $4\times$
@@
NeSR [44] 0.565 34.116 0.374 0.681 0.563 33.920 0.309 0.677
NeSSR [45] 0.359 38.130 0.323 0.849 0.436 35.441 0.431 0.788
[…]
IV-D Continuous Spectral Super-Resolution
To validate the effectiveness of our continuous modeling and evaluate the generalization and flexibility of our method, we conduct experiments on continuous spectral reconstruction. We train the model using only a small number of uniformly downsampled spectral bands and test it on the full spectral resolution. In our experiments, we apply $2\times$ and $4\times$ downsampling for continuous spectral reconstruction. We set $d_{\text{modes}}=12$ in this experiment to avoid $d_{\text{modes}}>C//2+1$ for high downsampling scales. We compare our method against two other continuous methods: NeSR and NeSSR. Table IV summarizes the quantitative results. Our method achieves best performance in both settings, indicating strong generalization to sparsely-sampled spectral bands.
The good performance of RSNO stems from two factors: the neural operator backbone ensuring resolution-invariant spectral interpolation, and the ACP method guided by radiative spectral prior ensuring radiative-structured spectral shape from seen to unseen bands. These results highlight RSNO’s robustness and generalization ability, making it suitable for practical hyperspectral applications.
@@ IV-E Ablation Study
where $\alpha$ is the trade-off parameter, and we set $\alpha=0.5$ in the ablation experiment. The results are listed in Table V. The results indicate that both the radiative prior and the final refinement play critical roles in overall performance, thereby demonstrating the effectiveness of the ACP method.
@@ V Conclusion
In this article, we propose a novel physics-informed approach for continuous spectral reconstruction. By continuously reformulating the problem and embedding physical priors within a data-driven framework, our method yields more realistic reconstructions and effectively balances interpretability and adaptability. Experiments across diverse real-world scenes and problem settings indicate better reconstruction accuracy and robustness, especially in the continuous setting. This work highlights the potential of continuous spectral modeling and the integration of domain knowledge with deep learning to advance spectral super-resolution for practical applications.